# Chinese Academy of Sciences, Beijing 100190, China

### v2

@@ Abstract
The primordial neutrino asymmetry leave profound imprints on the evolution history of the universe, which can be constrained by cosmological observations, including Big Bang Nucleosynthesis (BBN), Cosmic Microwave Background (CMB), and Large-Scale Structure (LSS). We present comprehensive analysis on implications and constraints of the primordial neutrino asymmetry $\xi_{\nu}$, based on a precise treatment of neutrino decoupling by solving the complete (anti)neutrino quantum kinetic equations in the Closed-Time-Path formalism. The effective number of neutrinos $N_{\rm eff}$ and (anti)neutrino spectral distortions are calculated, and we find that the non-instantaneous decoupling correction is given by $\delta N_{\rm eff}=0.0440+0.0102\,\xi_{\nu}^{2}$. Then we perform the state-of-the-art calculation for the abundance of light elements including (anti)neutrino spectral distortions, which indicate a positive asymmetry $0.032\leq\xi_{\nu}\leq 0.052$ from EMPRESS data. The implications of the neutrino asymmetry for the CMB and LSS are studied in detail, and we find that the Baryon Acoustic Oscillations (BAO) are also significantly affected by $\xi_{\nu}$ in addition to the sum of neutrino masses. A combined analysis with EMPRESS BBN, Planck CMB and BOSS BAO data yields a tighter constraint $\xi_{\nu}=0.024\pm 0.012$, which provides constraints on UV models capable of producing large asymmetries.
@@ 1 Introduction
[…]
@@
A precise understanding of the neutrino decoupling process is essential for accurate constraints on neutrino properties. Notably, the partial reheating of neutrinos and antineutrinos due to the small overlap between neutrino decoupling and $e^{\pm}$ annihilations results in slightly non-thermal spectra and an increased total neutrino energy density, typically described by the effective number of relativistic species, $N_{\rm eff}$. The values of $N_{\rm eff}$ and the spectral distortions of neutrinos and antineutrinos are used as initial conditions to evaluate the impact of relic neutrinos in subsequent processes. For the Standard Model case, where asymmetries between neutrinos and antineutrinos are assumed to be negligible, significant progress has been made in detailing the neutrino decoupling process over recent decades Hannestad:1995rs ; Dolgov:1997mb ; Mangano:2001iu ; Mangano:2005cc ; Grohs:2015tfy ; deSalas:2016ztq ; Akita:2020szl ; Froustey:2019owm ; EscuderoAbenza:2020cmq ; Froustey:2020mcq ; Bennett:2020zkv . Including flavor oscillations, matter effects, Finite-Temperature QED (FTQED) corrections up to $\mathcal{O}(3)$, and full neutrino-electron and neutrino-neutrino collision terms, the current standard prediction yields $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$ Froustey:2020mcq ; Bennett:2020zkv ; Akita:2020szl .
@@
On the other hand, inspired by the recent EMPRESS survey on primordial helium abundance empress and its potential to alleviate cosmological tensions Barenboim:2016lxv ; Yeung:2020zde ; Seto:2021tad ; Kumar:2022vee ; Yeung:2024krv , the scenario where asymmetries between neutrinos and antineutrinos are not negligible has recently attracted significant attention. Due to the sphaleron processes in the early universe Kuzmin:1985mm ; Khlebnikov:1988sr ; Harvey:1990qw ; Dreiner:1992vm , one might naively expect neutrino asymmetries, or lepton asymmetries, to be of the same order as the baryon asymmetry of the universe, which is strongly constrained by BBN and CMB observations: $\eta_{B}\equiv(n_{B}-n_{\bar{B}})/n_{\gamma}=(6.14\pm 0.04)\times 10^{-10}$ Planck:2018vyg , where $n_{B}$, $n_{\bar{B}}$, and $n_{\gamma}$ denote the number densities of baryons, antibaryons and photons, respectively. However, substantial lepton asymmetries compared to the baryon asymmetry before the neutrino decoupling epoch can still be generated by various models, see $e.g.,$ Casas:1997gx ; Dolgov:1989us ; Bajc:1997ky ; Asaka:2005pn ; Asaka:2005an ; Pilaftsis:2003gt ; Borah:2022uos ; Kawasaki:2002hq ; Kawasaki:2022hvx .
@@
The lepton asymmetries present before the neutrino decoupling epoch, henceforth referred to as primordial neutrino asymmetries, have significant implications for the universe’s evolution. During the neutrino decoupling epoch, these asymmetries contribute to an increased $N_{\rm eff}$ and induce specific spectral distortions in neutrinos and antineutrinos Grohs:2016cuu . If the asymmetries vary between different neutrino flavors, neutrinos and antineutrinos can undergo synchronous oscillations Bell:1998ds ; Pastor:2001iu ; Abazajian:2002qx ; Wong:2002fa , leading to a redistribution of asymmetries among the flavors Dolgov:2002ab ; Pastor:2008ti ; Mangano:2010ei ; Mangano:2011ip ; Castorina:2012md ; Barenboim:2016shh ; Johns:2016enc ; Froustey:2021azz . After neutrino decoupling, during the Big Bang Nucleosynthesis (BBN) epoch, primordial neutrino asymmetries primarily affect the final proton-to-neutron ratio and the resulting light element abundances. This effect depends largely on the distribution of electron neutrinos and antineutrinos, while corrections from the increased $N_{\rm eff}$ are secondary Sarkar:1995dd ; Iocco:2008va ; Pitrou:2018cgg ; Serpico:2005bc ; Mangano:2011ip ; Chu:2006ua ; Simha:2008mt ; Saviano:2013ktj ; Burns:2023sgx . Additionally, these asymmetries influence cosmic microwave background (CMB) anisotropies and large-scale structure (LSS) through the increased $N_{\rm eff}$ and modified light element abundances, along with the effects of massive neutrinos. Consequently, observations of BBN, CMB, and LSS can constrain primordial neutrino asymmetries Simha:2008mt ; Oldengott:2017tzj ; Pitrou:2018cgg ; Matsumoto:2022tlr ; Burns:2022hkq ; Escudero:2022okz . Notably, recent findings from the EMPRESS survey suggest a preference for a positive asymmetry in electron neutrinos Matsumoto:2022tlr ; Burns:2022hkq ; Escudero:2022okzx. Specifically, combining BBN and CMB observations yields $\xi_{\nu_{e}}=0.034\pm 0.014$ Escudero:2022okz , where $\xi_{\nu_{e}}=\mu_{\nu_{e}}/T_{\nu}$ is the degeneracy parameter for electron neutrinos.
@@
In this work, we focus on evaluating the neutrino decoupling process and its subsequent effects on BBN, CMB, and LSS in the presence of primordial neutrino asymmetries. To accurately describe the evolution of neutrino density matrices during decoupling, including flavor oscillations and full collision effects, we employ the complete set of neutrino quantum kinetic equations (QKEs). These equations are derived using several approaches, such as perturbative expansions of the density matrix Sigl:1993ctk , relativistic generalizations of the Bogoliubov-Born-Green-Kirkwood-Yvon (BBGKY) equations Volpe:2013uxl ; Volpe:2015rla ; Froustey:2020mcq , and the Closed-Time Path (CTP) formalism from non-equilibrium quantum field theory Vlasenko:2013fja ; Blaschke:2016xxt ; Kainulainen:2023ocv ; Kainulainen:2024fdg . The main improvements of our study are as follows:
@@
  • We numerically solve the QKEs for the neutrino decoupling process, accounting for flavor oscillations, matter interactions, FTQED corrections up to $\mathcal{O}(3)$, and complete neutrino-electron and neutrino-neutrino collision terms, matching the accuracy of state-of-the-art Standard Model (SM) predictions Froustey:2020mcq ; Bennett:2020zkv  [^1] Additional QED corrections, such as those to interaction vertices Cielo:2023bqp ; Jackson:2023zkl ; Drewes:2024wbw , remain under debate for the SM case and are not included in this study..
@@ 2.2 Gradient expansion of the kinetic equation
$\frac{\partial_{X},m,\Sigma,S^{H}}{E_{\nu}}=O(\epsilon)\,,\quad\frac{\tilde{\Pi}^{\pm}}{E_{\nu}}=O(\epsilon^{2})\,,$ (15)
@@
    Beyond $\mathcal{O}(\epsilon^{0})$, all spinor components of $S^{F}$ receive additional corrections due to neutrino masses and interactions. Thus, solve the kinetic equation (13b) at $\mathcal{O}(\epsilon^{1})$ to find the constraint relations among these small components.
@@
    Expand the kinetic equation (13b) to $\mathcal{O}(\epsilon^{2})$ and reorganize it using the constraint relations obtained at $\mathcal{O}(\epsilon^{1})$.
@@ 2.2.1 Kinetic equation to $\mathcal{O}(\epsilon^{0})$
Here, we have introduced a set of basis vectors used to express the Lorentz tensors and components of the two-point functions. The basis consists of two light-like vectors $\hat{\kappa}^{\mu}(k)=({\rm sign}(k^{0}),\hat{k})$ and $\hat{\kappa}^{\prime\mu}(k)=({\rm sign}(k^{0}),-\hat{k})$, and two transverse four-vectors $\hat{x}_{1,2}(k)$ which satisfy the relations $\hat{\kappa}\cdot\hat{\kappa}=\hat{\kappa}^{\prime}\cdot\hat{\kappa}^{\prime}=0$, $\hat{\kappa}\cdot\hat{\kappa}^{\prime}=2$, $\hat{\kappa}\cdot\hat{x}_{i}=\hat{\kappa}^{\prime}\cdot\hat{x}_{i}=0$, and $\hat{x}_{i}\cdot\hat{x}_{j}=-\delta_{ij}$. Additionally, we define $\hat{x}^{\pm}\equiv\hat{x}_{1}\pm i\hat{x}_{2}$, such that $\hat{x}^{+}\cdot\hat{x}^{-}=-2$.
@@
where $n_{f}$ is the total number of neutrino flavors, and each block is a square $n_{f}\times n_{f}$ matrix with $\varrho_{hh^{\prime}}=\varrho_{hh^{\prime}}^{ij}$, where $i,\,j$ denote the flavors and $h,\,h^{\prime}\in\{L,R\}$ denote the chirality. Among the various components, $\varrho_{LL}$ and $\bar{\varrho}_{RR}$ represent the density matrices for active left-handed neutrinos and right-handed antineutrinos, respectively. Conversely, $\varrho_{RR}$ and $\bar{\varrho}_{LL}$ correspond to sterile right-handed neutrinos and left-handed antineutrinos. The remaining density matrices account for the quantum coherence between different chiralities of neutrinos and antineutrinos across all flavors. Note that there is no single-particle density matrix for quantum coherence between Dirac neutrinos and antineutrinos, as they are fundamentally distinct particles.
@@
where $\varrho$ and $\bar{\varrho}$ denote the density matrices for active neutrinos and antineutrinos, respectively, and $\phi$ describes the quantum spin coherence between neutrinos and antineutrinos. Similar to the Dirac case, the density matrix for Majorana neutrinos, ${\cal F}$, can be extracted from the two-point functions via
$-2\int_{0}^{\infty}\frac{dk^{0}}{2\pi}\hat{F}(k,x)=\mathcal{F}(\vec{k},x)-\frac{1}{2}\mathbb{1}\,.$ (22)
@@
2.2.2 Kinetic equation to $\mathcal{O}(\epsilon^{1})$

### v3

# Primordial lepton asymmetries: neutrino transport, spectral distortions and cosmological constraints

@@ Abstract
The primordial neutrino asymmetry leaves profound imprints on the evolution history of the universe, which can be constrained by cosmological observations, including Big Bang Nucleosynthesis (BBN), Cosmic Microwave Background (CMB), and Large-Scale Structure (LSS). We present comprehensive analysis on implications and constraints of the primordial neutrino asymmetry, based on a precise treatment of neutrino decoupling by solving the complete (anti)neutrino quantum kinetic equations in the Closed-Time-Path formalism. Assuming the same primordial asymmetry $\xi_{\nu}$ for all neutrino flavors, the effective number of neutrinos $N_{\rm eff}$ and (anti)neutrino spectral distortions are calculated, and we find that the non-instantaneous decoupling correction is given by $\delta N_{\rm eff}=0.0440+0.0102\,\xi_{\nu}^{2}$. Then we perform the state-of-the-art calculation for the abundance of light elements including (anti)neutrino spectral distortions, which indicate a positive asymmetry $0.032\leq\xi_{\nu}\leq 0.052$ from EMPRESS data. The implications of the neutrino asymmetry for the CMB and LSS are studied in detail, and we find that the Baryon Acoustic Oscillations (BAO) are also significantly affected by $\xi_{\nu}$ in addition to the sum of neutrino masses. A combined analysis with EMPRESS BBN, Planck CMB and BOSS BAO data yields a tighter constraint $\xi_{\nu}=0.024\pm 0.012$, which provides constraints on UV models capable of producing large asymmetries.
@@ 1 Introduction
The presence of neutrinos, as evidenced by various cosmological observations, is crucial for understanding the evolution of the universe Lesgourgues:2013sjj . In the early universe, neutrinos and antineutrinos remain in thermal equilibrium with the Standard Model (SM) thermal bath as long as the interaction rates between (anti)neutrinos and other SM particles exceed the Hubble expansion rate. This equilibrium breaks down when the temperature falls below approximately $1\,\text{MeV}$, leading to the decoupling of neutrinos from the SM thermal bath. Shortly after decoupling, the annihilation of electrons and positrons reheats the SM thermal bath, causing the photon temperature to exceed that of the neutrinos, with $T_{\gamma}/T_{\nu}\simeq(11/4)^{1/3}$. Subsequent processes in the universe, in particular Big Bang Nucleosynthesis (BBN) and the formation of the Cosmic Microwave Background (CMB) anisotropies and Large Scale Structure (LSS), are both affected by the relic neutrinos, allowing constraints on neutrino properties to be obtained from relevant observations, see $e.g.,$ Pitrou:2018cgg ; Planck:2018vyg ; empress ; DESI:2024mwx .
@@
A precise understanding of the neutrino decoupling process is essential for accurate constraints on neutrino properties. Notably, the partial reheating of neutrinos and antineutrinos due to the small overlap between neutrino decoupling and $e^{\pm}$ annihilations results in slightly non-thermal spectra and an increased total neutrino energy density, typically described by the effective number of relativistic species, $N_{\rm eff}$. The values of $N_{\rm eff}$ and the spectral distortions of neutrinos and antineutrinos are then used as initial conditions to evaluate the impact of relic neutrinos in subsequent processes. For the Standard Model case, where asymmetries between neutrinos and antineutrinos are assumed to be negligible, significant progress has been made in detailing the neutrino decoupling process over recent decades Hannestad:1995rs ; Dolgov:1997mb ; Mangano:2001iu ; Mangano:2005cc ; Grohs:2015tfy ; deSalas:2016ztq ; Akita:2020szl ; Froustey:2019owm ; EscuderoAbenza:2020cmq ; Froustey:2020mcq ; Bennett:2020zkv . Including flavor oscillations, matter effects, Finite-Temperature QED (FTQED) corrections up to $\mathcal{O}(e^{3})$, where $e$ is the elementary electric charge, and full neutrino-electron and neutrino-neutrino collision terms, the current standard prediction yields $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$ Froustey:2020mcq ; Bennett:2020zkv ; Akita:2020szl .
On the other hand, inspired by the recent EMPRESS survey on primordial helium abundance empress and its potential to alleviate cosmological tensions Barenboim:2016lxv ; Yeung:2020zde ; Seto:2021tad ; Kumar:2022vee ; Yeung:2024krv , the scenario where asymmetries between neutrinos and antineutrinos are not negligible has recently attracted significant attention. Typically, such neutrino asymmetries can be parameterised with the degeneracy parameters $\xi_{\nu_{\alpha}}\equiv\mu_{\nu_{\alpha}}/T_{\nu_{\alpha}}$, defined as the chemical potential for the flavour $\alpha$ neutrino normalised to its temperature, so that
$\eta_{\nu}\equiv\frac{1}{n_{\gamma}}\sum_{\alpha=e,\mu,\tau}(n_{\nu_{\alpha}}-n_{\bar{\nu}_{\alpha}})=\frac{\pi^{2}}{33\zeta(3)}\sum_{\alpha=e,\mu,\tau}\left(\xi_{\nu_{\alpha}}+\frac{\xi_{\nu_{\alpha}}^{3}}{\pi^{2}}\right)\ ,$ (1)
@@
where $n_{\gamma}$ is the photon number density and $n_{\nu_{\alpha}}(n_{\bar{\nu}_{\alpha}})$ is the (anti)neutrino number density with flavour $\alpha$. For the second equality, the SM value for the neutrino-photon temperature ratio $T_{\nu_{i}}/T_{\gamma}=(4/11)^{1/3}$ is assumed. Beyond thermal equilibrium, an accurate $\eta_{\nu}$ must be obtained by a precise treatment of the neutrino decoupling process for given $\xi_{\nu_{\alpha}}$ as initial conditions. Due to the sphaleron processes in the early universe Kuzmin:1985mm ; Khlebnikov:1988sr ; Harvey:1990qw ; Dreiner:1992vm , one might naively expect neutrino asymmetries, or lepton asymmetries, to be of the same order as the baryon asymmetry of the universe, which is strongly constrained by BBN and CMB observations: $\eta_{B}\equiv(n_{B}-n_{\bar{B}})/n_{\gamma}=(6.14\pm 0.04)\times 10^{-10}$ Planck:2018vyg , where $n_{B}$ and $n_{\bar{B}}$ denote the number densities of baryons and antibaryons, respectively. However, substantial lepton asymmetries compared to the baryon asymmetry before the neutrino decoupling epoch can still be generated by various models, see $e.g.,$ Casas:1997gx ; Dolgov:1989us ; Bajc:1997ky ; Asaka:2005pn ; Asaka:2005an ; Pilaftsis:2003gt ; Borah:2022uos ; Kawasaki:2002hq ; Kawasaki:2022hvx .
@@
The lepton asymmetries present before the neutrino decoupling epoch, henceforth referred to as primordial neutrino asymmetries, have significant implications for the universe’s evolution. During the neutrino decoupling epoch, these asymmetries contribute to an increased $N_{\rm eff}$ and induce specific spectral distortions in neutrinos and antineutrinos Grohs:2016cuu . If the asymmetries vary between different neutrino flavors, neutrinos and antineutrinos can undergo synchronous oscillations Bell:1998ds ; Pastor:2001iu ; Abazajian:2002qx ; Wong:2002fa , leading to a redistribution of asymmetries among the flavors Dolgov:2002ab ; Pastor:2008ti ; Mangano:2010ei ; Mangano:2011ip ; Castorina:2012md ; Barenboim:2016shh ; Johns:2016enc ; Froustey:2021azz . After neutrino decoupling, during the Big Bang Nucleosynthesis (BBN) epoch, primordial neutrino asymmetries primarily affect the final proton-to-neutron ratio and the resulting light element abundances. This effect depends largely on the distribution of electron neutrinos and antineutrinos, while corrections from the increased $N_{\rm eff}$ are secondary Sarkar:1995dd ; Iocco:2008va ; Pitrou:2018cgg ; Serpico:2005bc ; Mangano:2011ip ; Chu:2006ua ; Simha:2008mt ; Saviano:2013ktj ; Burns:2023sgx . Additionally, these asymmetries influence cosmic microwave background (CMB) anisotropies and large-scale structure (LSS) through the increased $N_{\rm eff}$ and modified light element abundances, along with the effects of massive neutrinos. Consequently, observations of BBN, CMB, and LSS can constrain primordial neutrino asymmetries Simha:2008mt ; Oldengott:2017tzj ; Pitrou:2018cgg ; Matsumoto:2022tlr ; Burns:2022hkq ; Escudero:2022okz . Notably, recent findings from the EMPRESS survey suggest a preference for a positive asymmetry in electron neutrinos Matsumoto:2022tlr ; Burns:2022hkq ; Escudero:2022okz . Specifically, combining BBN and CMB observations yields $\xi_{\nu_{e}}=0.034\pm 0.014$ Escudero:2022okz .
@@
In this work, we focus on evaluating the neutrino decoupling process and its subsequent effects on BBN, CMB, and LSS in the presence of primordial neutrino asymmetries. Since the synchronous oscillations before neutrino decoupling will redistribute the neutrino asymmetries between different flavors, the resulting neutrino asymmetries of different flavors are usually of the same order, so that $\xi_{\nu_{e}}\simeq\xi_{\nu_{\mu}}\simeq\xi_{\nu_{\tau}}$ Dolgov:2002ab ; Pastor:2008ti ; Mangano:2010ei ; Mangano:2011ip ; Castorina:2012md ; Barenboim:2016shh ; Johns:2016enc ; Froustey:2021azz . Therefore, a complete flavor equilibration with $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$ is usually assumed in the study of the cosmological constraints on neutrino asymmetries Simha:2008mt ; Oldengott:2017tzj ; Pitrou:2018cgg ; Matsumoto:2022tlr ; Burns:2022hkq ; Escudero:2022okz . The framework we have developed in this work can also be straightforwardly applied to studies of the general flavour equilibration process. However, a detailed illustration is beyond the scope of this paper. Therefore, we will also assume complete flavor equilibration in this paper and leave further studies of general flavor equilibration to future work. Relevant previous discussions of the exact flavor equilibration process and its cosmological implications can also be found in, $e.g.,$ Pastor:2008ti ; Mangano:2011ip ; Froustey:2024mgf . To accurately describe the evolution of neutrino density matrices during decoupling, including flavor oscillations and full collision effects, we employ the complete set of neutrino quantum kinetic equations (QKEs). These equations are derived using several approaches, such as perturbative expansions of the density matrix Sigl:1993ctk , relativistic generalizations of the Bogoliubov-Born-Green-Kirkwood-Yvon (BBGKY) equations Volpe:2013uxl ; Volpe:2015rla ; Froustey:2020mcq , and the Closed-Time Path (CTP) formalism from non-equilibrium quantum field theory Vlasenko:2013fja ; Blaschke:2016xxt ; Kainulainen:2023ocv ; Kainulainen:2024fdg . The main improvements of our study are as follows:
@@
  • We numerically solve the QKEs for the neutrino decoupling process, accounting for flavor oscillations, matter interactions, FTQED corrections up to $\mathcal{O}(e^{3})$, and complete neutrino-electron and neutrino-neutrino collision terms, matching the accuracy of state-of-the-art Standard Model (SM) predictions Froustey:2020mcq ; Bennett:2020zkv  [^1] Additional QED corrections, such as those to interaction vertices Cielo:2023bqp ; Jackson:2023zkl ; Drewes:2024wbw , remain under debate for the SM case and are not included in this study..
@@ 2.2 Gradient expansion of the kinetic equation
$\frac{\partial_{X},m,\Sigma,S^{H}}{E_{\nu}}=\mathcal{O}(\epsilon)\,,\quad\frac{\tilde{\Pi}^{\pm}}{E_{\nu}}=\mathcal{O}(\epsilon^{2})\,$, (16)
@@
    Beyond $\mathcal{O}(\epsilon^{0})$, all spinor components of $S^{F}$ receive additional corrections due to neutrino masses and interactions. Thus, solve the kinetic equation (14b) at $\mathcal{O}(\epsilon)$ to find the constraint relations among these small components.
@@
    Expand the kinetic equation (14b) to $\mathcal{O}(\epsilon^{2})$ and reorganize it using the constraint relations obtained at $\mathcal{O}(\epsilon)$.
@@ 2.2.1 Kinetic equation to $\mathcal{O}(\epsilon^{0})$
[…]
@@
where $n_{f}$ is the total number of neutrino flavors, and each block is a square $n_{f}\times n_{f}$ matrix with $\varrho_{hh^{\prime}}=\varrho_{hh^{\prime}}^{ij}$, where i,ji,\,j denote the flavors and $h,\,h^{\prime}\in\{L,R\}$ denote the chirality. Among the various components, $\varrho_{LL}$ and $\bar{\varrho}_{RR}$ represent the density matrices for active left-handed neutrinos and right-handed antineutrinos, respectively. Conversely, $\varrho_{RR}$ and $\bar{\varrho}_{LL}$ correspond to sterile right-handed neutrinos and left-handed antineutrinos. The remaining density matrices account for the quantum coherence between different chiralities of neutrinos and antineutrinos across all flavors. Note that there is no single-particle density matrix for quantum coherence between Dirac neutrinos and antineutrinos, as they are fundamentally distinct particles.
@@
[…]
$2\int_{0}^{\infty}\frac{dk^{0}}{2\pi}\hat{F}(k,x)=\mathcal{F}(\vec{k},x)-\frac{1}{2}\mathbb{1}\,.$ (23)
@@
2.2.2 Kinetic equation to $\mathcal{O}(\epsilon)$